OpenStudy (anonymous):

Evaluate the expression when x = 6. 3x + 7 A. 25 B. 39 C. 43 D. 126

OpenStudy (acxbox22):

3(6)+7=18+7=25

OpenStudy (solomonzelman):

plug in 6 for x